The Smith family and the Cooper family each used their sprinklers last summer. The Smith family's sprinkler was used for 15 hours. The Cooper family's
sprinkler was used for 25 hours. There was a combined total output of 1075 L of water. What was the water output rate for each sprinkler if the sum of the two
rates was 55 L per hour?

Answer by Boreal(15235)

x L/h Smith so 15x L
55-x L/h Cooper so 25(55-x) L
they add to 1075 L
15x+1375-25x=1075
-10x=-300
x=30 L/hr Smith for 450 L
55-x=25 L/hr Cooper for 625 L